# Towards an Understanding of Context Utilization in Code Intelligence

### v1

@@ Abstract.
ccs: Software and its engineering Automatic programming
@@ 1. Introduction
Code intelligence (CI) is a rapidly growing field that aims to harness the potential of artificial intelligence techniques, particularly deep learning, to extract and utilize knowledge embedded in large-scale code repositories(Xu and Zhu, 2022; Sun et al., 2024a). This domain supports critical software engineering tasks such as code generation (Athiwaratkun et al., 2023; Zheng et al., 2025b), program repair (Li et al., 2020; Wen et al., 2018), and code summarization (Yu et al., 2020; Shi et al., 2021), all of which are crucial for improving software development, maintenance, and evolution (Wang et al., 2024h; Zheng et al., 2025a). Recent advancements in the field have focused on incorporating contextual information, which provides additional information beyond the basic inputs required for these tasks. For example, CoCoGen (Bi et al., 2024b) leverages compiler feedback to iteratively localize project-level context and refine the generated code, ensuring that project-specific details are preserved. Similarly, Li et al. (Li, 2020) propose a context-based bug detection method that reduces false positives by integrating three types of context: program dependence graphs (PDGs), data flow graphs (DFGs), and abstract syntax trees (ASTs). Furthermore, RLPG (Shrivastava et al., 2023b) takes the repository structure into account and trains a prompt proposal classifier to effectively utilize repository context for code completion, and RepoCoder (Zhang et al., 2023a) proposes an iterative retrieval-generation pipeline to collect relevant code from the repository as additional context.
@@
  • RQ2: How are contexts categorized in code intelligence tasks? We investigate different types of context used in code intelligence and form a taxonomy. Based on whether contexts can be obtained directly, we categorize them into two major groups: direct context (e.g., source code, code diffs, API documents) and indirect context (e.g., ASTs, control flow graphs). Analysis results show that direct context is more frequently used than indirect context. Among the 14 identified forms of context, the most commonly used are source code, abstract syntax trees (ASTs), and code diffs. Although context is widely leveraged, there are still specific areas where the application is limited. The extent of context utilization varies significantly across tasks depending on how well the context aligns with the specific requirements of each task, revealing underexplored opportunities in cross-context generalization.
@@
The remainder of this paper is organized as follows: Section 2 introduces the background of code intelligence tasks. Section 3 describes the methodology used for paper selection. Sections IV to VII address RQs, respectively. Section 8 presents related work and research focuses. Section 9 provides a detailed explanation to current challenges and potential opportunities. Section 10 concludes our study and outlines directions for future research in context utilization.
@@ 2. Background
Code intelligence, as defined by Wan et al. (Wan et al., 2024), refers to the application of artificial intelligence techniques to extract knowledge from code repositories and assist programmers in various code-related tasks (Shi et al., 2023, 2022; Zheng et al., 2024b; Tao et al., 2021; Wang et al., 2024e, c; Gong et al., 2024; Guo et al., 2024b). Code intelligence is pivotal in software engineering, equipping developers with advanced tools that reduce manual effort and boost productivity. In this subsection, we briefly introduce seven key code intelligence tasks examined in this paper, including four code-code tasks (clone detection, defect detection, code completion, program repair), two code-text tasks (code summarization, commit message generation), and a text-code task (code generation).
@@
  • Code Generation: Code generation, also known as program synthesis, involves automatically generating source code from natural language descriptions. Early approaches were limited to predefined scenarios (Mattingley et al., 2010) or required manual template construction (Charfi et al., 2010), resulting in poor generalizability. The advent of machine learning (Ling et al., 2016; Yang et al., 2021; Rabinovich et al., 2017) has enabled models to generate code for diverse domains, alleviating template dependency. However, unlike human developers who intrinsically utilize context (e.g., compiler information, API documents), early ML models lacked mechanisms to incorporate such knowledge. Recent advances address this gap by integrating contextual cues. For example, API documentation (Zhou et al., 2023; Washio and Miyao, 2022) for semantic constraints, codebase analysis (Liao et al., 2024; Wang et al., 2024g) for structural patterns, and compiler feedback (Bi et al., 2024b; Wang et al., 2022) for syntactic validation. These context-aware techniques significantly improve code relevance and accuracy, demonstrating the potential of repository-grounded code generation.
@@
  • Code Completion: Code completion aims to generate syntactically correct and semantically coherent code fragments for unfinished code snippets. This task can be categorized along two orthogonal dimensions: (1) the type of uncompleted code, including subsequent code prediction (Lu et al., 2022) and code infilling (Bavarian et al., 2022); (2) the granularity of completion, ranging from token-level (Zhu et al., 2024) and line-level (Naznin and Mondal, 2023) to function-level (Zhang et al., 2023a) completion. Effective context utilization is pivotal across all types of code completion, as local variables, cross-file dependencies, and API calls necessitate a deep understanding of the development context. This requirement becomes particularly critical in repository-level completion (Zhang et al., 2023a; Phan et al., 2024; Wu et al., 2024b; Li et al., 2024e), where a retrieval module collects project-specific contexts to maintain global code consistency. Recent advances adopt two complementary strategies to enhance repository-level code completion: (1) Structural approaches like context graph construction (Ding et al., 2023b; Phan et al., 2024) or code dependencies extraction to optimize retrieval methods; (2) Pipeline-centric methods employ iterative retrieval (Zhang et al., 2023a), dynamic judgement (Wu et al., 2024b; Tan et al., 2024), and reranking (Li et al., 2024e) to search for efficient contexts and filter noisy context progressively. These innovations collectively advance the precision of context-aware code completion.
@@
  • Code Summarization: Code summarization, known as code comment generation, aims to automatically produce concise and informative descriptions based on given source code. This task significantly enhances software maintainability and team collaboration (Zhang et al., 2022; Wang et al., 2024b). Early approaches rely on rigid rule-based templates (Abid et al., 2015), and recent advances leverage deep learning to improve semantic understanding: CNNs extract local code features (Allamanis et al., 2016), RNNs model sequential logic (Iyer et al., 2016), GNNs capture syntactic structures (Zhang et al., 2021), and Transformers address long-range dependencies (Guo et al., 2024a). However, even state-of-the-art architectures struggle to infer implicit knowledge (e.g., API usage conventions) from standalone code snippets. To bridge this gap, researchers enrich models with various types of context such as source code (Zhou et al., 2019) for syntactic patterns, API documents (Shahbazi and Fard, 2023) for semantic constraints, code comments (Wong et al., 2015) for intent clarification, and UML diagrams (Wang et al., 2021a) for architectural context. By jointly analyzing repositories, context-aware models generate summaries that better articulate functional core and systemic roles, advancing precision and comprehensibility.
@@
  • Commit Message Generation: The commit message generation task takes code diffs as essential input and aims to automatically produce natural language summaries that concisely describe what code changes were made and why they were implemented (Tian et al., 2022; Tao et al., 2024). High-quality commit messages are critical for software maintenance (Zhang et al., 2024c), as they document repository evolution and facilitate collaboration. Early rule-based methods (Buse and Weimer, 2010) extract keywords from source code to compose messages but suffer from redundant text and fail to explain change rationales (Zhang et al., 2024d). Approaches based on deep learning models (Liu et al., 2019; Loyola et al., 2018, 2017)) improved fluency by learning from commit histories, yet often generate vague messages (e.g., ”optimize performance”) due to ignoring project-specific context. This context blindness causes style inconsistency and functional ambiguity. To address this, researchers first explore retrieving messages from similar historical commits (Huang et al., 2017), but retrieval alone struggles with novel code changes. Recent hybrid methods (Wang et al., 2021b; Li et al., 2024a; Imani et al., 2024) integrate neural generators with retrieval mechanisms: retrievers fetch context to guide style alignment, while generators synthesize commit messages with current changes and retrieved contexts. Such methods demonstrate promise in balancing specificity and consistency, advancing practical applicability.
@@
  • Defect Detection: Defect detection identifies code fragments that may introduce functional errors, security vulnerabilities, or performance degradation. Its predictive capability significantly enhances software reliability (Thota et al., 2020), driving active research across academia and industry. Current studies target diverse defect categories, such as just-in-time defects (Zhao et al., 2023), reentry vulnerabilities (Liu et al., 2018), and security-critical flaws from the Common Vulnerabilities and Exposures (CVE) (Wang et al., 2024a). While deep learning models (Li, 2020; Zhang et al., 2023b; Liu et al., 2022) and code similarity analysis (Sun et al., 2021; Cui et al., 2021) achieve notable progress, they still exhibit critical limitations without the guidance of context. For instance, without modeling cross-file dependencies in code property graphs, models may misclassify secure code as vulnerable when interacting with external libraries (Liu et al., 2022). To address this, recent works emphasize models with multi-level contextual information, such as bug reports (Youm et al., 2015), code change history (Youm et al., 2017)), and context graph representations (Zhang et al., 2023b; Liu et al., 2022). Thereby enables more precise analysis in defect detection.
@@
  • Program Repair: Program repair task involves taking buggy source code as input and producing a corrected version while maintaining the original functionality. Existing approaches span rule-based methods using heuristic search (Qi et al., 2014; Sun et al., 2018), manual restoration using templates (Tian and Ray, 2017; Liu and Zhong, 2018), and the application of semantic constraints (Mechtaev et al., 2015; Lee et al., 2018). The rise of deep learning also opens new research directions (Abhinav et al., 2021; Li et al., 2020), as deep neural networks can automatically extract features, potentially freeing researchers from labor-intensive rule writing and feature engineering. Automatic program repair remains challenging due to the complexity of handling interdependent code fragments and external libraries, which increases the need for contextual information to guide repairs. Incorporating contextual information from the source code (Lutellier et al., 2020; Nashid et al., 2023; Kim et al., 2019) helps narrow down potential bug sources, enabling more accurate identification and resolution of issues. Additionally, compiler information (Li et al., 2023c) provides valuable context regarding syntax and execution, highlighting areas where code fails to compile or execute correctly. Despite the exciting progress in automatic program repair, a significant gap remains in achieving repository-level automatic program repair (Chen et al., 2024b), which demands further exploration in context utilization and model improvement.
@@ 3. Methodology
To gain a comprehensive understanding of context utilization in the code intelligence domain, we focus on seven key tasks: four code-to-code tasks (clone detection, defect detection, code completion, and program repair), two code-to-text tasks (code summarization and commit message generation), and one text-to-code task (code generation).
@@
[…]
[Short description]The Pipeline in Collecting Relevant Papers
@@ 3.1. Research Questions
These research questions are designed to address key gaps in understanding what kind of context is utilized and how context is applied across various code intelligence tasks. RQ1 focuses on the historical trends and distribution of publications, offering insights into the growing interest in context-aware models across seven key code intelligence tasks since 2007. RQ2 classifies different context types and analyzes context’s impact across different scenarios. RQ3 explores the preprocessing methods and techniques presented in recent studies to optimize context usage and analyzes the strengths and weaknesses of these approaches. Finally, RQ4 investigates the evaluation of context-aware models, including the metrics and datasets employed and their open-source status, to identify potential areas for improvement in model evaluation.
By answering these research questions, we aim to provide a comprehensive overview of context utilization in code intelligence and offer guidelines for selecting and applying the appropriate techniques based on specific contexts.
@@ 3.2. Literature Search and Selection
We follow the methodology illustrated in Figure 1 to collect papers on using context in code intelligence. Initially, we employ an open card sorting method (Chen et al., 2021b) to define two categories: keywords and tasks. To broaden the scope of the search, we also include aliases for code intelligence tasks, such as code comment generation and vulnerability detection. We then search Google Scholar using combinations of keywords and tasks, as shown in Table 1. Each keyword is paired with each task, and we conduct an exhaustive search until all possible combinations are explored. For papers that meet our inclusion/exclusion criteria (outlined in Secton 3.3), we apply a snowballing strategy (Zheng et al., 2024a) for each associated paper, reviewing both the references cited by this paper and the papers that cite this paper. As a result, we identify 86 pertinent papers in the preliminary investigation. After reviewing these papers, we categorize the types of context, as detailed in Table 1. In the comprehensive stage, we repeatedly pair each context with each task for further searches, applying the snowballing strategy again. Ultimately, we collected 163 papers and retained 146 after confirming their quality.
@@
[…]
[Short description]Words Used for Searching

### v2

@@ Abstract.
$c$s: Software and its engineering Automatic programming
@@ 1. Introduction
Code intelligence (CI) is a rapidly growing field that aims to harness the potential of artificial intelligence techniques, particularly deep learning, to extract and utilize knowledge embedded in large-scale code repositories(Xu and Zhu, 2022; Sun et al., 2024). This domain supports critical software engineering tasks such as code generation (Athiwaratkun et al., 2023), program repair (Li et al., 2020), and code summarization (Yu et al., 2020), all of which are crucial for improving software development, maintenance, and evolution (Wang et al., 2024h; Zheng et al., 2025b). Recent advancements in the field have focused on incorporating contextual information, which provides additional information beyond the basic inputs required for these tasks. For example, CoCoGen (Bi et al., 2024) leverages compiler feedback to iteratively localize project-level context and refine the generated code, ensuring that project-specific details are preserved. Similarly, Li et al. (Li, 2020) propose a context-based bug detection method that reduces false positives by integrating three types of context: program dependence graphs (PDGs), data flow graphs (DFGs), and abstract syntax trees (ASTs). Furthermore, RLPG (Shrivastava et al., 2023b) takes the repository structure into account and trains a prompt proposal classifier to effectively utilize repository context for code completion, and RepoCoder (Zhang et al., 2023a) proposes an iterative retrieval-generation pipeline to collect relevant code from the repository as additional context.
@@
  • RQ2: How are contexts categorized in code intelligence tasks? We investigate different types of context used in code intelligence and form a taxonomy. Based on whether contexts can be obtained directly, we categorize them into two major groups: direct context (e.g., source code, code diffs, API documents) and indirect context (e.g., ASTs, control flow graphs). Analysis results show that direct context is more frequently used than indirect context. Among the 14 identified forms of context, the most commonly used are source code, abstract syntax trees (ASTs), and code diffs. Although context is widely leveraged, there are still specific areas where the application is limited. The extent of context utilization varies significantly across tasks depending on how well the context aligns with the specific requirements of each task, revealing underexplored opportunities in context generalization across multiple tasks and the effective utilization of multiple context types within a single task.
@@
Besides, we maintain a GitHub project [^1] https://github.com/DeepSoftwareAnalytics/Repo-Context-Survey associated with this survey, which contains the resources used to construct this paper. We will continue to update the resource, hoping that this approach will benefit the community and facilitate future research.
The remainder of this paper is organized as follows: Section 2 introduces the background of code intelligence tasks. Section 3 describes the methodology used for paper selection. Sections 4 to 7 address RQs, respectively. Section 8 presents a case study that compares performance with the addition of context. Section 9 presents related work and research focuses. Section 10 provides a detailed explanation of current challenges and potential opportunities. Section 11 concludes our study and outlines directions for future research in context utilization.
@@ 2. Background
Code intelligence, as defined by Wan et al. (Wan et al., 2024), refers to the application of artificial intelligence techniques to extract knowledge from code repositories and assist programmers in various code-related tasks (Shi et al., 2023, 2022; Zheng et al., 2024a; Tao et al., 2021; Wang et al., 2024d, f; Gong et al., 2024; Guo et al., 2024a). Code intelligence is pivotal in software engineering, equipping developers with advanced tools that reduce manual effort and boost productivity. Recent advances suggest that incorporating contextual information is the key to achieving better performance in code intelligence tasks. Context is the supplementary information beyond the raw input (the minimal, essential data a task requires), that a model leverages to enhance its comprehension and effectiveness. In this subsection, we briefly introduce seven key code intelligence tasks examined in this paper, including four code-code tasks (clone detection, defect detection, code completion, program repair), two code-text tasks (code summarization, commit message generation), and a text-code task (code generation). The specific definition and raw input for each task are detailed in Table 1.
@@
  • Code Generation: Code generation, also known as program synthesis, involves automatically generating source code from natural language descriptions. Early approaches were limited to predefined scenarios (Mattingley et al., 2010) or required manual template construction (Charfi et al., 2010), resulting in poor generalizability. The advent of machine learning (Ling et al., 2016; Yang et al., 2021; Rabinovich et al., 2017) has enabled models to generate code for diverse domains, alleviating template dependency. However, unlike human developers who intrinsically utilize context (e.g., compiler information, API documents), early ML models lacked mechanisms to incorporate such knowledge. Recent advances address this gap by integrating contextual cues. For example, API documentation (Zhou et al., 2023; Washio and Miyao, 2022) for semantic constraints, codebase analysis (Liao et al., 2024; Wang et al., 2024a) for structural patterns, and compiler feedback (Bi et al., 2024; Wang et al., 2022) for syntactic validation. These context-aware techniques significantly improve code relevance and accuracy, demonstrating the potential of repository-grounded code generation.
@@
  • Code Completion: Code completion aims to generate syntactically correct and semantically coherent code fragments for unfinished code snippets. This task can be categorized along two orthogonal dimensions: (1) the type of uncompleted code, including subsequent code prediction (Lu et al., 2022) and code infilling (Bavarian et al., 2022); (2) the granularity of completion, ranging from token-level (Zhu et al., 2024) and line-level (Naznin and Mondal, 2023) to function-level (Zhang et al., 2023a) completion. Effective context utilization is pivotal across all types of code completion, as local variables, cross-file dependencies, and API calls necessitate a deep understanding of the development context. This requirement becomes particularly critical in repository-level completion (Zhang et al., 2023a; Phan et al., 2024; Wu et al., 2024a; Li et al., 2024f), where a retrieval module collects project-specific contexts to maintain global code consistency. Recent advances adopt two complementary strategies to enhance repository-level code completion: (1) Structural approaches like context graph construction (Ding et al., 2023b; Phan et al., 2024) or code dependencies extraction to optimize retrieval methods; (2) Pipeline-centric methods employ iterative retrieval (Zhang et al., 2023a), dynamic judgement (Wu et al., 2024a; Tan et al., 2024), and reranking (Li et al., 2024f) to search for efficient contexts and filter noisy context progressively. These innovations collectively advance the precision of context-aware code completion.
@@
  • Code Summarization: Code summarization, known as code comment generation, aims to automatically produce concise and informative descriptions based on given source code. This task significantly enhances software maintainability and team collaboration (Zhang et al., 2022; Wang et al., 2024e). Early approaches rely on rigid rule-based templates (Abid et al., 2015), and recent advances leverage deep learning to improve semantic understanding: CNNs extract local code features (Allamanis et al., 2016), RNNs model sequential logic (Iyer et al., 2016), GNNs capture syntactic structures (Zhang et al., 2021), and Transformers address long-range dependencies (Guo et al., 2024b). However, even state-of-the-art architectures struggle to infer implicit knowledge (e.g., API usage conventions) from standalone code snippets. To bridge this gap, researchers enrich models with various types of context such as source code (Zhou et al., 2019) for syntactic patterns, API documents (Shahbazi and Fard, 2023) for semantic constraints, code comments (Wong et al., 2015) for intent clarification, and UML diagrams (Wang et al., 2021b) for architectural context. By jointly analyzing repositories, context-aware models generate summaries that better articulate functional core and systemic roles, advancing precision and comprehensibility.
@@
  • Commit Message Generation: The commit message generation task takes code diffs as essential input and aims to automatically produce natural language summaries that concisely describe what code changes were made and why they were implemented (Tian et al., 2022; Tao et al., 2024). High-quality commit messages are critical for software maintenance (Zhang et al., 2024f), as they document repository evolution and facilitate collaboration. Early rule-based methods (Buse and Weimer, 2010) extract keywords from source code to compose messages but suffer from redundant text and fail to explain change rationales (Zhang et al., 2024e). Approaches based on deep learning models (Liu et al., 2019; Loyola et al., 2018, 2017)) improved fluency by learning from commit histories, yet often generate vague messages (e.g., ”optimize performance”) due to ignoring project-specific context. This context blindness causes style inconsistency and functional ambiguity. To address this, researchers first explore retrieving messages from similar historical commits (Huang et al., 2017), but retrieval alone struggles with novel code changes. Recent hybrid methods (Wang et al., 2021a; Li et al., 2024d; Imani et al., 2024) integrate neural generators with retrieval mechanisms: retrievers fetch context to guide style alignment, while generators synthesize commit messages with current changes and retrieved contexts. Such methods demonstrate promise in balancing specificity and consistency, advancing practical applicability.
@@
  • Defect Detection: Defect detection identifies code fragments that may introduce functional errors, security vulnerabilities, or performance degradation. Its predictive capability significantly enhances software reliability (Thota et al., 2020), driving active research across academia and industry. Current studies target diverse defect categories, such as just-in-time defects (Zhao et al., 2023), reentry vulnerabilities (Liu et al., 2018), and security-critical flaws from the Common Vulnerabilities and Exposures (CVE) (Wang et al., 2024c). While deep learning models (Li, 2020; Zhang et al., 2023b; Liu et al., 2022) and code similarity analysis (Sun et al., 2021; Cui et al., 2021) achieve notable progress, they still exhibit critical limitations without the guidance of context. For instance, without modeling cross-file dependencies in code property graphs, models may misclassify secure code as vulnerable when interacting with external libraries (Liu et al., 2022). To address this, recent works emphasize models with multi-level contextual information, such as bug reports (Youm et al., 2015), code change history (Youm et al., 2017)), and context graph representations (Zhang et al., 2023b; Liu et al., 2022). This enables more precise analysis in defect detection.
@@
  • Program Repair: Program repair task involves taking buggy source code as input and producing a corrected version while maintaining the original functionality. Existing approaches span rule-based methods using heuristic search (Qi et al., 2014; Sun et al., 2018), manual restoration using templates (Tian and Ray, 2017; Liu and Zhong, 2018), and the application of semantic constraints (Mechtaev et al., 2015; Lee et al., 2018). The rise of deep learning also opens new research directions (Abhinav et al., 2021; Li et al., 2020), as deep neural networks can automatically extract features, potentially freeing researchers from labor-intensive rule writing and feature engineering. Automatic program repair remains challenging due to the complexity of handling interdependent code fragments and external libraries, which increases the need for contextual information to guide repairs. Incorporating contextual information from the source code (Lutellier et al., 2020; Nashid et al., 2023; Kim et al., 2019) helps narrow down potential bug sources, enabling more accurate identification and resolution of issues. Additionally, compiler information (Li et al., 2023b) provides valuable context regarding syntax and execution, highlighting areas where code fails to compile or execute correctly. Despite the exciting progress in automatic program repair, a significant gap remains in achieving repository-level automatic program repair (Chen et al., 2024), which demands further exploration in context utilization and model improvement.
Task Definition Raw Input
Code Generation Generate corresponding source code based on natural language descriptions. Natural language descriptions
Code Completion Predict code elements (e.g., tokens, lines, or blocks) based on the current incomplete code. Incomplete code units
Code Summarization Generate concise and descriptive comments based on the given source code. Source code
Program Repair Automatically fix defects (bugs) in the given source code. Buggy source code
Clone Detection Identify code snippets with notable similarities. Code snippet pairs
Defect Detection Identify whether the given code snippet contains potential bugs or vulnerabilities. Source code snippets
Commit Message Generation Automatically generate descriptive commit messages for code changes (diffs). Code diffs
Table 1. the Definition and the Raw Input of Each Code Intelligence Task.
@@ 3. Methodology
To gain a comprehensive understanding of context utilization in the code intelligence domain, we follow previous work (Sun et al., 2024) and select seven key tasks based on their input/output modalities. This structure ensures representative coverage across the three fundamental categories: four code-to-code tasks (clone detection, defect detection, code completion, and program repair), two code-to-text tasks (code summarization and commit message generation), and one text-to-code task (code generation).
@@
Short description The Pipeline in Collecting Relevant Papers
@@ 3.1. Research Questions
These research questions are designed to address key gaps in understanding which context is used and how it is applied across various code intelligence tasks. RQ1 examines historical trends and the distribution of publications, offering insights into the growing interest in context-aware models across seven key code intelligence tasks since 2007. RQ2 classifies context types and analyzes their impact across scenarios. RQ3 explores the preprocessing methods and techniques presented in recent studies to optimize context usage and analyzes the strengths and weaknesses of these approaches. Finally, RQ4 evaluates context-aware models, including the metrics and datasets used and their open-source status, to identify potential areas for improvement in model evaluation.
By answering these research questions, we aim to provide a comprehensive overview of context utilization in code intelligence and offer guidelines for selecting and applying techniques appropriate to specific contexts.
@@ 3.2. Literature Search and Selection
We follow the methodology illustrated in Figure 1 to collect papers on using context in code intelligence. We employ a systematic, multi-phase process to define our search query and ensure completeness. Our search is divided into two primary stages: a preliminary investigation using initial keywords to broadly identify relevant work, followed by a comprehensive stage that uses context types gathered from the preliminary set to refine and complete the search. We identified 86 pertinent papers in the preliminary investigation and collected 163 papers after the comprehensive stage. After confirming their quality, we retained 146 papers.
Search Query Definition. Initially, we define two categories: Tasks (Specific, actionable research problems) and Keywords (Broad, high-level terms defining the research domain). To ensure representative coverage, the set of seven code intelligence tasks is pre-confirmed by structuring our selection based on three fundamental input/output modalities. To ensure maximal comprehensive search scope, we explicitly include aliases for code intelligence tasks (e.g., “code summarization” and “code comment generation”; “defect detection” and “vulnerability detection”) as distinct entries in the tasks list. During the preliminary investigation, we employ an open card-sorting method (Chen et al., 2021a), in which five authors leverage their domain expertise to generate keywords (broad conceptual terms, such as “context” and “repository”) and consolidate the initial term pool. During the comprehensive stage, we categorize the identified context types after reviewing the preliminary papers. These context types replace keywords for further searches. The details of keywords, context types, and tasks are shown in Table 2.
Literature Retrieval and Completeness Assurance. Three rigorous methods are employed to ensure the completeness of our literature set.
  • Phased Investigation: Our literature search is executed in two distinct stages: a preliminary investigation and a comprehensive stage. Initially, the preliminary investigation searches 3,240 papers primarily using Google Scholar  [^2] https://scholar.google.com/. For the comprehensive stage, we expand the search data sources and utilize the context types identified in the preliminary investigation as refined keywords, resulting in a larger search scope of 22,680 papers.
  • Iterative Exhaustion: In the preliminary investigation, we iteratively use the “AND” operator to combine the Keyword and Task for searching, until all combinations are explored. After reviewing the preliminary papers, we categorized the identified context types, as detailed in Table 1. In the comprehensive stage, we repeatedly paired each context type (acting as new refined keywords) with each task for further searches.
  • Two-Stage Snowballing: For papers that meet our inclusion/exclusion criteria (outlined in Secton 3.3), we apply a snowballing strategy (Zheng et al., 2024c) in both investigation stages (reviewing both the references cited by this paper and the papers that cite this paper).
@@
Short description Words Used for Searching